OpenStudy (anonymous):

Janise rolled a number cube 50 times and recorded the results in this table. Outcome Number 1 7 2 9 3 8 4 8 5 6 6 12 What is the theoretical probability of the cube landing on a 6? What is the experimental probability of the cube landing on a 6? A. Theoretical - 16.7% Experimental - 16.7% B. Theoretical - 16.7% Experimental - 24% C. Theoretical - 20% Experimental - 12% D. Theoretical - 12% Experimental - 24%

OpenStudy (anonymous):

@CO_oLBoY

OpenStudy (anonymous):

@ranga @ganeshie8 @Mel98

OpenStudy (anonymous):

@parakeet

OpenStudy (anonymous):

theoretical = (1/6) x 100 % =16.7% experimental= (12/50)x100 % =24%

OpenStudy (anonymous):

So B

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

see i gave u your medal

OpenStudy (anonymous):

A coin is tossed 10 times resulting in 7 heads and 3 tails. The same coin is tossed 1000 times resulting in 510 heads and 490 tails. What is the theoretical probability of getting heads with this coin? A. 51% B. 49% C. 70% D. 50%

OpenStudy (anonymous):

i think it is B

OpenStudy (anonymous):

its D... the ques says theoretical prob. which is (1/2) x 100=50%

OpenStudy (anonymous):

are u sure

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Larry drew a blue marble from a bag and, without replacing it, drew a red marble from the same bag. Are these events dependent or independent? A. dependent B. independent

OpenStudy (anonymous):

ask every que in a new que so u can give medals to every ans ;)

OpenStudy (anonymous):

okay i will post it wait

OpenStudy (anonymous):

btw the ans is A.dependent